40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4493     Entry time: Wed Apr 6 18:55:49 2011
Author: Jamie, Larisa 
Type: Configuration 
Category: LSC 
Subject: major AP table cleanup 

We ripped out all of the old AS, PLL, and REFL paths, green, orange, and cyan respectively on the old AP table layout photo:

  • AS (green): had already been re-purposed by putting a ThorLabs diode right after the first steering mirror.   Everything downstream of that has been removed.
  • PLL (orange): everything removed.
  • REFL (cyan): CCD was left in place, so everything upstream of that was not touched.  Everything else was removed, including all of the REFL detectors.
  • OMCT (purple): previously removed
  • OMCR (blue): left in place, but the diode and CCD are not connected (found that way).
  • MCT (magenta): previously removed.
  • IMRC (red): untouched

All optics and components were moved to the very south end of the SP table.

We also removed all spurious cables from the table top, and from underneath, as well as pulled out no-longer-needed power supplies.

ELOG V3.1.3-